GNU LibreJS

for version 7.21, 17 September 2023

Loic J. Duros (librejs@lduros.net)
Yuchen Pei (ycp@gnu.org)

This manual is for GNU LibreJS (version 7.21, 17 September 2023), a GNU IceCat extension to detect and block nonfree nontrivial JavaScript on webpages. Copyright \begin{matrix}c \\ ◯ \end{matrix} 2011 2012 2014 2015 Loic J. Duros, 2022 Yuchen Pei

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, with no Front-Cover Texts, and with no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Table of Contents

1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Disclaimer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Installation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

4 How to Use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4.1 LibreJS in action . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.2 Script Blacklist/whitelist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.3 Complaint Feature. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.4 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

5 JavaScript Detection . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

6 Free Licenses Detection . . . . . . . . . . . . . . . . . . . . . . . . . 7

7 Setting Your JavaScript Free . . . . . . . . . . . . . . . . . . . 8

7.1 License tags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 7.2 Undetected Free Licenses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 7.3 Known limitations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

8 LibreJS Development Notes. . . . . . . . . . . . . . . . . . . 19

8.1 Dependencies. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.2 Building. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.3 Debugging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.4 Testing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.5 Headless testing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.6 Headless compliance check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 8.7 Adding new whitelisted libraries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 8.8 Releasing a new version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

Appendix A Installation Requirements . . . . . . . . . 22

A.1 Mozilla Browser. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

Appendix B LibreJS Internals. . . . . . . . . . . . . . . . . . . 23

Appendix C GNU Free Documentation License . . 24

1 Overview

GNU LibreJS —an add-on for GNU IceCat and Mozilla Firefox— detects and blocks nonfree nontrivial JavaScript while allowing its execution on pages containing code that is either trivial and/or free.

Many websites run nontrivial JavaScript on your computer. Some use it for complex tasks; many use it gratuitously for minor jobs that could be done easily with plain HTML. Sometimes this JavaScript code is malicious. Either way, the JavaScript code is often nonfree. For explanation of the issue, see "The JavaScript Trap"(http://www.gnu.org/

philosophy/javascript-trap.html).

If you care about freedom in your computing, and don’t wish to let all and sundry make you run nonfree programs, now you can prevent it by using LibreJS.

2 Disclaimer

• LibreJS is not a security tool. Its goal is to detect nonfree nontrivial JavaScript, and it currently does not detect whether free or trivial code is malicious or not. Other free Mozilla extensions and add-ons may be available for this purpose.

• LibreJS is always a work in progress. If you find a bug, please report it to

bug-librejs@gnu.org.

3 Installation

You can install LibreJS directly using a generated librejs.xpi file, or by building it from source (See Chapter 8 [LibreJS Development Notes], page 19).

   You can also download it from https://www.gnu.org/software/librejs/ or https://

addons.mozilla.org/addon/librejs/, but due to Mozilla’s review process the download isn’t always up to date.

4 How to Use

4.1 LibreJS in action

After installing the add-on, you will see the LibreJS widget in the add-on bar at the top right of the browser window. After loading a page, left-click on the widget to view the deactivated JavaScript code from the page (both on page and external) and, if applicable, the scripts that were accepted.

4.2 Script Blacklist/whitelist

Scripts may be blacklisted or whitelisted through the pop-up menu box.

Whitelisted scripts will always be accepted and blacklisted scripts will always get rejected.

It is important to note that this feature recognizes which scripts are blacklisted and whitelisted based on hash. This means that even a slight difference in a script’s code will cause it to be recognized as a separate script.

Sometimes, JavaScript will be dynamically generated so that it is different every time a website is loaded. These types of scripts cannot be whitelisted or blacklisted since they cannot be recognized.

LibreJS has a default whitelist of scripts that are known to be free but may not declare their license in a format that LibreJS can understand.

4.3 Complaint Feature

It is very important to complain when a site has nonfree JavaScript code, especially if it won’t work without that code. LibreJS makes it easy to complain by heuristically finding where to send the complaint.

When nonfree/nontrivial code is detected in a page, LibreJS attempts to find a relevant contact link or email for the website you are visiting. In order to do so, it will attempt to visit a few links from the current page (for instance, a link labeled “contact” on the same domain as the current page, …)

LibreJS detects contact pages, email addresses that are likely to be owned by the maintainer of the site, Twitter and identi.ca links, and phone numbers.

When you complain to the website for their nonfree nontrivial JavaScript, provide them with the link to the JavaScript Trap essay so that they can get more information on what the issue is and how they can solve it on their own site.

LibreJS includes a default subject line and body for the complaint email, with a link to the JavaScript Trap essay. This can be configured in the LibreJS add-on preferences in your web browser.

4.4 Options

You can manage LibreJS’s preferences either from the extension’s entry in your browser’s Add-ons Manager page (about:addons) or by clicking the LibreJS toolbar icon and then the "Settings..." button on the top right of the popup. This will open a panel containing a whitelist/blacklist manager and a section to configure your complaints messages to site owners.

Whitelist/Blacklist LibreJS lets you whitelist or blacklist domain names and subdomains, to bypass the regular JavaScript checks. This might be useful, for example, if you are running your own code in a local web server, or if you don’t want to waste computing resources on script origins you already know you can’t trust. librejs provides a lists manager UI to handle both the lists on the top of its Options panel.

Complaint email subject Configure the default subject used in complaint emails.

Complaint email body Configure the default body used in complaint emails.

5 JavaScript Detection

LibreJS considers a very strict subset of JavaScript to be acceptable for use in non-free scripts. This is meant to maximimize compatibility with websites that haven’t tried to be LibreJS compatible.

We consider modification of the document non-trivial. There isn’t much that javascript could do that we would consider trivial, for anything else a free software license would be required.

The criterion is as follows: For each function definition:

• It must call only primitives.

• The number of conditionals and loops must be at most 3.

• It does not declare an array more than 50 elements long.

• It must not call itself For the rest of the script, outside of function definitions:

• It must call only primitives and functions defined above in the page.

• The number of conditionals and loops must be at most 3. "function" means anything executable that gets a name, including methods. Allowed primitives exclude:

• eval()

• ajax

• calling methods with the square bracket notation

• altering the dom

• most other items found as methods of the ‘.window‘ object.

6 Free Licenses Detection

The machine readable format for license declarations that LibreJS uses has changed in the most recent version. This was necessary in order to not break the asynchronous JS loading model that browsers use. Scripts are now evaluated independent of eachother and strictly as they arrive.

7 Setting Your JavaScript Free

The first step is releasing your JavaScript under a free license. If you are already using a free library, or you’re not using any third-party libraries, it might only take a few minutes.

On your website, take a look at your HTML source. You can identify distinct pieces of JavaScript that might be free and some other that are nonfree.

This might be the case with an analytics tracker, social media widgets, and code that runs ads. Removing these pieces of code from your site is required to have the rest accepted as free. There are often alternatives to nonfree libraries or to third-party services:

• If you have used nonfree third-party code as the base to write your own code, try to find a free alternative.

• If you’re using a third-party service such as an analytics service, replace it with a free alternative like Matomo.

• If you can’t find free JavaScript that has already been developed, write it yourself! Who knows, your own solution might be the start of a brilliant project!

7.1 License tags

LibreJS will allow non-trivial scripts to run as long as they use a free license.

In order for the license of a script to be recognized by LibreJS, it must be declared using a machine-readable license format.

This format is the same for both remote in-line scripts.

"// @license [magnet link] [identifier]" [Script here] "// @license-end"

"Identifier" is a name of a license from the following list and the magnet link in the canonicalUrl field is that license’s exact corresponding magnet link.

The following json object which can be found in ./common/license_definitions.json file in the LibreJS code shows all the licensees recognised by LibreJS.

{
  "AGPL-3.0": {
    "licenseName": "GNU AFFERO GENERAL PUBLIC LICENSE version 3",
    "identifier": "AGPL-3.0",
    "canonicalUrl": [
      "http://www.gnu.org/licenses/agpl-3.0.html",
      "magnet:?xt=urn:btih:0b31508aeb0634b347b8270c7bee4d411b5d4109&dn=agpl-3.0.txt"
    ],
    "licenseFragments": [
      {
        "text": "<THISPROGRAM> is free software: you can redistribute it and/or modify it under the terms of the GNU Affero General Public License as published by the Free Software Foundation, either version 3 of the License, or (at your option) any later version.",
        "type": "short"
      }
    ]
  },
  "Apache-2.0": {
    "licenseName": "Apache License, Version 2.0",
    "identifier": "Apache-2.0",

"canonicalUrl": [
  "http://www.apache.org/licenses/LICENSE-2.0",
  "magnet:?xt=urn:btih:8e4f440f4c65981c5bf93c76d35135ba5064d8b7&dn=apache-2.0.txt"
],
"licenseFragments": [
  {
    "text": "Licensed under the Apache License, Version 2.0 (the \"License\"); you may not use this file except in compliance with the License. You may obtain a copy of the License at http://www.apache.org/licenses/LICENSE-2.0",
    "type": "short"
  }
]
},
"Artistic-2.0": {
"licenseName": "Artistic License 2.0",
"identifier": "Artistic-2.0",
"canonicalUrl": [
  "http://www.perlfoundation.org/artistic_license_2_0",
  "magnet:?xt=urn:btih:54fd2283f9dbdf29466d2df1a98bf8f65cafe314&dn=artistic-2.0.txt"
],
"licenseFragments": []
},
"BSD-2-Clause": {
"licenseName": "BSD 2-Clause License",
"identifier": "BSD-2-Clause",
"canonicalUrl": [
  "http://www.freebsd.org/copyright/freebsd-license.html",
  "magnet:?xt=urn:btih:87f119ba0b429ba17a44b4bffcab33165ebdacc0&dn=freebsd.txt"
],
"licenseFragments": [
  {
    "text": "Redistribution and use in source and binary forms, with or without modification, are permitted provided that the following conditions are met: Redistributions of source code must retain the above copyright notice, this list of conditions and the following disclaimer. Redistributions in binary form must reproduce the above copyright notice, this list of conditions and the following disclaimer in the documentation and/or other materials provided with the distribution.",
    "type": "short"
  }
]
},
"BSD-3-Clause": {
"licenseName": "BSD 3-Clause License",
"identifier": "BSD-3-Clause",
"canonicalUrl": [
  "http://opensource.org/licenses/BSD-3-Clause",
  "magnet:?xt=urn:btih:c80d50af7d3db9be66a4d0a86db0286e4fd33292&dn=bsd-3-clause.txt"
],
"licenseFragments": [
  {
    "text": "Redistribution and use in source and binary forms, with or without modification, are permitted provided that the following conditions are met: Redistributions of source code must retain the above copyright notice, this list of conditions and the following disclaimer. Redistributions in binary form must reproduce the above copyright notice, this list of conditions and the following disclaimer in the documentation and/or other materials provided with the distribution. Neither the name of <ORGANIZATION> nor the names of its contributors may be used to endorse or promote products derived from this software without specific prior written permission.",
    "type": "short"
  }
]

},
"BSL-1.0": {
"licenseName": "Boost Software License 1.0",
"identifier": "BSL-1.0",
"canonicalUrl": [
  "http://www.boost.org/LICENSE_1_0.txt",
  "magnet:?xt=urn:btih:89a97c535628232f2f3888c2b7b8ffd4c078cec0&dn=Boost-1.0.txt"
],
"licenseFragments": [
  {
    "text": "Boost Software License <VERSION> <DATE> Permission is hereby granted, free of charge, to any person or organization obtaining a copy of the software and accompanying documentation covered by this license (the \"Software\") to use, reproduce, display, distribute, execute, and transmit the Software, and to prepare derivative works of the Software, and to permit third-parties to whom the Software is furnished to do so, all subject to the following",
    "type": "short"
  }
]
},
"CC-BY-1.0": {
"licenseName": "Creative Commons Attribution 1.0 Generic",
"identifier": "CC-BY-1.0",
"canonicalUrl": [
  "https://creativecommons.org/licenses/by/1.0/"
],
"licenseFragments": []
},
"CC-BY-2.0": {
"licenseName": "Creative Commons Attribution 2.0 Generic",
"identifier": "CC-BY-2.0",
"canonicalUrl": [
  "https://creativecommons.org/licenses/by/2.0/"
],
"licenseFragments": []
},
"CC-BY-2.5": {
"licenseName": "Creative Commons Attribution 2.5 Generic",
"identifier": "CC-BY-2.5",
"canonicalUrl": [
  "https://creativecommons.org/licenses/by/2.5/"
],
"licenseFragments": []
},
"CC-BY-3.0": {
"licenseName": "Creative Commons Attribution 3.0 Unported",
"identifier": "CC-BY-3.0",
"canonicalUrl": [
  "https://creativecommons.org/licenses/by/3.0/"
],
"licenseFragments": []
},

"CC-BY-4.0": {
"licenseName": "Creative Commons Attribution 4.0 International",
"identifier": "CC-BY-4.0",
"canonicalUrl": [
  "https://creativecommons.org/licenses/by/4.0/"
],
"licenseFragments": []
},
"CC-BY-SA-1.0": {
"licenseName": "Creative Commons Attribution-ShareAlike 1.0 Generic",
"identifier": "CC-BY-SA-1.0",
"canonicalUrl": [
  "https://creativecommons.org/licenses/by-sa/1.0/"
],
"licenseFragments": []
},
"CC-BY-SA-2.0": {
"licenseName": "Creative Commons Attribution-ShareAlike 2.0 Generic",
"identifier": "CC-BY-SA-2.0",
"canonicalUrl": [
  "https://creativecommons.org/licenses/by-sa/2.0/"
],
"licenseFragments": []
},
"CC-BY-SA-2.5": {
"licenseName": "Creative Commons Attribution-ShareAlike 2.5 Generic",
"identifier": "CC-BY-SA-2.5",
"canonicalUrl": [
  "https://creativecommons.org/licenses/by-sa/2.5/"
],
"licenseFragments": []
},
"CC-BY-SA-3.0": {
"licenseName": "Creative Commons Attribution-ShareAlike 3.0 Unported",
"identifier": "CC-BY-SA-3.0",
"canonicalUrl": [
  "https://creativecommons.org/licenses/by-sa/3.0/"
],
"licenseFragments": []
},
"CC-BY-SA-4.0": {
"licenseName": "Creative Commons Attribution-ShareAlike 4.0 International",
"identifier": "CC-BY-SA-4.0",
"canonicalUrl": [
  "https://creativecommons.org/licenses/by-sa/4.0/"
],
"licenseFragments": []

},
"CC0-1.0": {
"licenseName": "Creative Commons CC0 1.0 Universal",
"identifier": "CC0-1.0",
"canonicalUrl": [
  "http://creativecommons.org/publicdomain/zero/1.0/legalcode",
  "magnet:?xt=urn:btih:90dc5c0be029de84e523b9b3922520e79e0e6f08&dn=cc0.txt"
],
"licenseFragments": []
},
"CECILL-2.0": {
"licenseName": "CeCILL Free Software License Agreement v2.0",
"identifier": "CECILL-2.0",
"canonicalUrl": [
  "https://www.cecill.info/licences/Licence_CeCILL_V2-en.txt",
  "magnet:?xt=urn:btih:dda0473d240d7febeac8fa265da27286ead0b1ce&dn=cecill-2.0.txt"
],
"licenseFragments": []
},
"CPAL-1.0": {
"licenseName": "Common Public Attribution License Version 1.0 (CPAL)",
"identifier": "CPAL-1.0",
"canonicalUrl": [
  "http://opensource.org/licenses/cpal_1.0",
  "magnet:?xt=urn:btih:84143bc45939fc8fa42921d619a95462c2031c5c&dn=cpal-1.0.txt"
],
"licenseFragments": [
  {
    "text": "The contents of this file are subject to the Common Public Attribution License Version 1.0",
    "type": "short"
  },
  {
    "text": "The term \"External Deployment\" means the use, distribution, or communication of the Original Code or Modifications in any way such that the Original Code or Modifications may be used by anyone other than You, whether those works are distributed or communicated to those persons or made available as an application intended for use over a network. As an express condition for the grants of license hereunder, You must treat any External Deployment by You of the Original Code or Modifications as a distribution under section 3.1 and make Source Code available under Section 3.2.",
    "type": "short"
  }
]
},
"EPL-1.0": {
"licenseName": "Eclipse Public License Version 1.0",
"identifier": "EPL-1.0",
"canonicalUrl": [
  "http://www.eclipse.org/legal/epl-v10.html",
  "magnet:?xt=urn:btih:4c6a2ad0018cd461e9b0fc44e1b340d2c1828b22&dn=epl-1.0.txt"
],
"licenseFragments": [
  {
    "text": "THE ACCOMPANYING PROGRAM IS PROVIDED UNDER THE TERMS OF THIS ECLIPSE PUBLIC LICENSE (\"AGREEMENT\"). ANY USE, REPRODUCTION OR DISTRIBUTION OF THE PROGRAM CONSTITUTES RECIPIENT'S ACCEPTANCE OF THIS AGREEMENT.",

    "type": "short"
  }
]
},
"Expat": {
"licenseName": "Expat License (sometimes called MIT Licensed)",
"identifier": "Expat",
"canonicalUrl": [
  "http://www.jclark.com/xml/copying.txt",
  "magnet:?xt=urn:btih:d3d9a9a6595521f9666a5e94cc830dab83b65699&dn=expat.txt"
],
"licenseFragments": [
  {
    "text": "Copyright <YEAR> <NAME> Permission is hereby granted, free of charge, to any person obtaining a copy of this software and associated documentation files (the \"Software\"), to deal in the Software without restriction, including without limitation the rights to use, copy, modify, merge, publish, distribute, sublicense, and/or sell copies of the Software, and to permit persons to whom the Software is furnished to do so, subject to the following conditions: The above copyright notice and this permission notice shall be included in all copies or substantial portions of the Software.",
    "type": "short"
  }
]
},
"FreeBSD": {
"licenseName": "FreeBSD License",
"identifier": "FreeBSD",
"canonicalUrl": [
  "http://www.freebsd.org/copyright/freebsd-license.html",
  "magnet:?xt=urn:btih:87f119ba0b429ba17a44b4bffcab33165ebdacc0&dn=freebsd.txt"
],
"licenseFragments": [
  {
    "text": "Redistribution and use in source and binary forms, with or without modification, are permitted provided that the following conditions are met:\n\nRedistributions of source code must retain the above copyright notice, this list of conditions and the following disclaimer.\n\nRedistributions in binary form must reproduce the above copyright notice, this list of conditions and the following disclaimer in the documentation and/or other materials provided with the distribution.",
    "type": "short"
  }
]
},
"GNU-All-Permissive": {
"licenseName": "GNU All-Permissive License",
"identifier": "GNU-All-Permissive",
"canonicalUrl": [],
"licenseFragments": [
  {
    "text": "Copying and distribution of this file, with or without modification, are permitted in any medium without royalty provided the copyright notice and this notice are preserved. This file is offered as-is, without any warranty.",
    "type": "short"
  }
]
},
"GPL-2.0": {
"licenseName": "GNU General Public License (GPL) version 2",
"identifier": "GPL-2.0",
"canonicalUrl": [

"http://www.gnu.org/licenses/gpl-2.0.html",
"magnet:?xt=urn:btih:cf05388f2679ee054f2beb29a391d25f4e673ac3&dn=gpl-2.0.txt"
],
"licenseFragments": [
{
  "text": "<THISPROGRAM> is free software; you can redistribute it and/or modify it under the terms of the GNU General Public License as published by the Free Software Foundation; either version 2 of the License, or (at your option) any later version.",
  "type": "short"
},
{
  "text": "Alternatively, the contents of this file may be used under the terms of either the GNU General Public License Version 2 or later (the \"GPL\"), or the GNU Lesser General Public License Version 2.1 or later (the \"LGPL\"), in which case the provisions of the GPL or the LGPL are applicable instead of those above. If you wish to allow use of your version of this file only under the terms of either the GPL or the LGPL, and not to allow others to use your version of this file under the terms of the MPL, indicate your decision by deleting the provisions above and replace them with the notice and other provisions required by the GPL or the LGPL. If you do not delete the provisions above, a recipient may use your version of this file under the terms of any one of the MPL, the GPL or the LGPL.",
  "type": "short"
}
]
},
"GPL-3.0": {
"licenseName": "GNU General Public License (GPL) version 3",
"identifier": "GPL-3.0",
"canonicalUrl": [
"http://www.gnu.org/licenses/gpl-3.0.html",
"magnet:?xt=urn:btih:1f739d935676111cfff4b4693e3816e664797050&dn=gpl-3.0.txt"
],
"licenseFragments": [
{
  "text": "<THISPROGRAM> is free software: you can redistribute it and/or modify it under the terms of the GNU  General Public License (GNU GPL) as published by the Free Software  Foundation, either version 3 of the License, or (at your option)  any later version. The code is distributed WITHOUT ANY WARRANTY; without even the implied warranty of MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE. See the GNU GPL for more details. As additional permission under GNU GPL version 3 section 7, you may distribute non-source (e.g., minimized or compacted) forms of that code without the copy of the GNU GPL normally required by section 4, provided you include this license notice and a URL through which recipients can access the Corresponding Source.",
  "type": "short"
},
{
  "text": "<THISPROGRAM> is free software: you can redistribute it and/or modify it under the terms of the GNU General Public License as published by the Free Software Foundation, either version 3 of the License, or (at your option) any later version.",
  "type": "short"
}
]
},
"ISC": {
"licenseName": "The ISC License",
"identifier": "ISC",
"canonicalUrl": [
"https://www.isc.org/downloads/software-support-policy/isc-license/",
"magnet:?xt=urn:btih:b8999bbaf509c08d127678643c515b9ab0836bae&dn=ISC.txt"
],
"licenseFragments": [
{
  "text": "Permission to use, copy, modify, and/or distribute this software for any purpose with or without fee is hereby granted, provided that the above copyright notice and this permission notice appear in all copies.\n\nTHE SOFTWARE IS PROVIDED \"AS IS\" AND ISC DISCLAIMS ALL WARRANTIES WITH REGARD TO THIS SOFTWARE INCLUDING ALL IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS. IN NO EVENT SHALL ISC BE LIABLE FOR ANY SPECIAL, DIRECT, INDIRECT, OR CONSEQUENTIAL DAMAGES OR ANY DAMAGES WHATSOEVER RESULTING FROM LOSS OF USE, DATA OR PROFITS, WHETHER IN AN ACTION OF CONTRACT, NEGLIGENCE OR OTHER TORTIOUS ACTION, ARISING OUT OF OR IN CONNECTION WITH THE USE OR PERFORMANCE OF THIS SOFTWARE.",
  "type": "short"
},
{
  "text": "Permission to use, copy, modify, and/or distribute this software for any purpose with or without fee is hereby granted, provided that the above copyright notice and this permission notice appear in all copies.THE SOFTWARE IS PROVIDED \"AS IS\" AND THE AUTHOR DISCLAIMS ALL WARRANTIES WITH REGARD TO THIS SOFTWARE INCLUDING ALL IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS. IN NO EVENT SHALL THE AUTHOR BE LIABLE FOR ANY SPECIAL, DIRECT, INDIRECT, OR CONSEQUENTIAL DAMAGES OR ANY DAMAGES WHATSOEVER RESULTING FROM LOSS OF USE, DATA OR PROFITS, WHETHER IN AN ACTION OF CONTRACT, NEGLIGENCE OR OTHER TORTIOUS ACTION, ARISING OUT OF OR IN CONNECTION WITH THE USE OR PERFORMANCE OF THIS SOFTWARE.",
  "type": "short"

  }
]
},
"jQueryTools": {
"licenseName": "jQuery Tools",
"identifier": "jQueryTools",
"canonicalUrl": [],
"licenseFragments": [
  {
    "text": "NO COPYRIGHTS OR LICENSES. DO WHAT YOU LIKE.",
    "type": "short"
  }
]
},
"LGPL-2.1": {
"licenseName": "GNU Lesser General Public License, version 2.1",
"identifier": "LGPL-2.1",
"canonicalUrl": [
  "http://www.gnu.org/licenses/lgpl-2.1.html",
  "magnet:?xt=urn:btih:5de60da917303dbfad4f93fb1b985ced5a89eac2&dn=lgpl-2.1.txt"
],
"licenseFragments": [
  {
    "text": "<THISLIBRARY> is free software; you can redistribute it and/or modify it under the terms of the GNU Lesser General Public License as published by the Free Software Foundation; either version 2.1 of the License, or (at your option) any later version.",
    "type": "short"
  }
]
},
"LGPL-3.0": {
"licenseName": "GNU Lesser General Public License, version 3",
"identifier": "LGPL-3.0",
"canonicalUrl": [
  "http://www.gnu.org/licenses/lgpl-3.0.html",
  "magnet:?xt=urn:btih:0ef1b8170b3b615170ff270def6427c317705f85&dn=lgpl-3.0.txt"
],
"licenseFragments": [
  {
    "text": "<THISPROGRAM> is free software: you can redistribute it and/or modify it under the terms of the GNU Lesser General Public License as published by the Free Software Foundation, either version 3 of the License, or (at your option) any later version.",
    "type": "short"
  }
]
},
"MPL-2.0": {
"licenseName": "Mozilla Public License Version 2.0",
"identifier": "MPL-2.0",
"canonicalUrl": [
  "http://www.mozilla.org/MPL/2.0",

  "magnet:?xt=urn:btih:3877d6d54b3accd4bc32f8a48bf32ebc0901502a&dn=mpl-2.0.txt"
],
"licenseFragments": [
  {
    "text": "This Source Code Form is subject to the terms of the Mozilla Public License, v. 2.0. If a copy of the MPL was not distributed with this file, You can obtain one at http://mozilla.org/MPL/2.0/.",
    "type": "short"
  }
]
},
"PublicDomain": {
"licenseName": "Public Domain",
"identifier": "PublicDomain",
"canonicalUrl": [
  "magnet:?xt=urn:btih:e95b018ef3580986a04669f1b5879592219e2a7a&dn=public-domain.txt"
],
"licenseFragments": []
},
"Unlicense": {
"licenseName": "Unlicense",
"identifier": "Unlicense",
"canonicalUrl": [
  "http://unlicense.org/UNLICENSE",
  "magnet:?xt=urn:btih:5ac446d35272cc2e4e85e4325b146d0b7ca8f50c&dn=unlicense.txt"
],
"licenseFragments": [
  {
    "text": "This is free and unencumbered software released into the public domain.",
    "type": "short"
  }
]
},
"UPL": {
"licenseName": "Universal Permissive License",
"identifier": "UPL-1.0",
"canonicalUrl": [
  "https://oss.oracle.com/licenses/upl/",
  "magnet:?xt=urn:btih:478974f4d41c3fa84c4befba25f283527fad107d&dn=upl-1.0.txt"
],
"licenseFragments": [
  {
    "text": "The Universal Permissive License (UPL), Version 1.0",
    "type": "short"
  }
]
},
"WTFPL": {
"licenseName": "Do What The F*ck You Want To Public License (WTFPL)",

"identifier": "WTFPL",
"canonicalUrl": [
  "http://www.wtfpl.net/txt/copying/",
  "magnet:?xt=urn:btih:723febf9f6185544f57f0660a41489c7d6b4931b&dn=wtfpl.txt"
],
"licenseFragments": [
  {
    "text": "DO WHAT THE FUCK YOU WANT TO PUBLIC LICENSE",
    "type": "short"
  },
  {
    "text": "0. You just DO WHAT THE FUCK YOU WANT TO.",
    "type": "short"
  }
]
},
"X11": {
"licenseName": "X11 License",
"identifier": "X11",
"canonicalUrl": [
  "magnet:?xt=urn:btih:5305d91886084f776adcf57509a648432709a7c7&dn=x11.txt"
],
"licenseFragments": [
  {
    "text": "Permission is hereby granted, free of charge, to any person obtaining a copy of this software and associated documentation files (the \"Software\"), to deal in the Software without restriction, including without limitation the rights to use, copy, modify, merge, publish, distribute, sublicense, and/or sell copies of the Software, and to permit persons to whom the Software is furnished to do so, subject to the following conditions: The above copyright notice and this permission notice shall be included in all copies or substantial portions of the Software.",
    "type": "short"
  }
]
},
"XFree86-1.1": {
"licenseName": "XFree86 1.1 License",
"identifier": "XFree86-1.1",
"canonicalUrl": [
  "http://www.xfree86.org/3.3.6/COPYRIGHT2.html#3",
  "http://www.xfree86.org/current/LICENSE4.html",
  "magnet:?xt=urn:btih:12f2ec9e8de2a3b0002a33d518d6010cc8ab2ae9&dn=xfree86.txt"
],
"licenseFragments": [
  {
    "text": "All rights reserved.\nPermission is hereby granted, free of charge, to any person obtaining a copy of this software and associated documentation files (the \"Software\"), to deal in the Software without restriction, including without limitation the rights to use, copy, modify, merge, publish, distribute, sublicense, and/or sell copies of the Software, and to permit persons to whom the Software is furnished to do so, subject to the following conditions:\n1. Redistributions of source code must retain the above copyright notice, this list of conditions, and the following disclaimer.\n2. Redistributions in binary form must reproduce the above copyright notice, this list of conditions and the following disclaimer in the documentation and/or other materials provided with the distribution, and in the same place and form as other copyright, license and disclaimer information.\n3. The end-user documentation included with the redistribution, if any, must include the following acknowledgment: \"This product includes software developed by The XFree86 Project, Inc (http://www.xfree86.org/) and its contributors\", in the same place and form as other third-party acknowledgments. Alternately, this acknowledgment may appear in the software itself, in the same form and location as other such third-party acknowledgments.4. Except as contained in this notice, the name of The XFree86 Project, Inc shall not be used in advertising or otherwise to promote the sale, use or other dealings in this Software without prior written authorization from The XFree86 Project, Inc.",
    "type": "short"
  }
]
},
"Zlib": {
"licenseName": "zlib License",
"canonicalUrl": [

      "https://www.zlib.net/zlib_license.html",
      "https://spdx.org/licenses/Zlib.txt",
      "magnet:?xt=urn:btih:922bd98043fa3daf4f9417e3e8fec8406b1961a3&dn=zlib.txt"
    ],
    "identifier": "Zlib",
    "licenseFragments": []
  }
}

7.2 Undetected Free Licenses

If you are using a free license that isn’t detected by LibreJS and isn’t listed in the previous section, please send a message to bug-librejs@gnu.org regarding this license, where code released under this license can be found, and where to find the license text and information.

Many free licenses are listed in this page: http://www.gnu.org/licenses/

license-list.html

7.3 Known limitations

Service workers may cause false positives and false negatives, and there are multiple ways to disable them. The cleanest way is by setting dom.serviceWorkers.enabled to false in the about:config page of your browser.

8 LibreJS Development Notes

8.1 Dependencies

LibreJS 7.21 depends on a number of Node.js-based libraries that can be installed using the npm utility:

  $ npm install acorn@"<=8.7.1" jssha browserify
  $ export PATH=$PATH:./node_modules/.bin

8.2 Building

To build the extension run:

  $ browserify main_background.js -o bundle.js

To build the extension plus create a .xpi package run:

  $ ./build.sh

To build the extension including the automated test suite (see TEST below) run:

  $ ./build.sh -t

or

  $ ./build.sh --test

Note: this build.sh script relies on no new source files being created.

8.3 Debugging

To debug LibreJS, visit the special URL about:debugging. Click on ‘Enable add-on debugging‘ then ‘Load Temporary Add-on‘. Navigate to LibreJS’s unpacked source directory

and select manifest.json.

Lines 39 and 40 in main_background.js assign two variables controlling the verbosity of dbg_print() statements. Make sure these are set to false before building a release.

8.4 Testing

An automated test suite runs automatically in its own tab whenever the extension is loaded as a "Temporary add-on" from about:debugging. Otherwise (if included in the xpi) it can be launched from the UI by clicking the "Automated self test..." button.

8.5 Headless testing

To launch the test suite from the command line, ensure the xpi package has been built with automated test suite. Then install selenium-webdriver and geckodriver, and ensure the latter is in $PATH:

  $ npm install selenium-webdriver geckodriver
  $ export PATH=$PATH:./node_modules/.bin

Now you can invoke the test with

  $ node ./utilities/test.js

which will print out a summary of test results.

Optionally you can also test with a chosen seed

 $ node ./utilities/test.js 12345

8.6 Headless compliance check

To check whether a webpage is LibreJS-compliant from the command line, ensure that the xpi package has been built WITHOUT the automated test suite. Then install seleniumwebdriver and geckodriver, and ensure the latter is in $PATH:

 $ npm install selenium-webdriver
 $ npm install geckodriver
 $ export PATH=$PATH:./node_modules/.bin

Now you can check a webpage for compliance with

 $ node ./utilities/compliance.js <url>

It will open the url in a headless browser, save a screenshot, and output the compliance check result.

For example, to check the compliance of the FSF homepage, do

 $ node ./utilities/compliance.js https://fsf.org

8.7 Adding new whitelisted libraries

The script index.js in ./utilities/hash_script generates the default whitelist. Run it with the following command:

  node index.js > output

Then, just copy the contents of the file "output" to the appropriate place in main background.js.

8.8 Releasing a new version

1. Run some tests and make sure they all pass.

2. Make sure debug statements are set to false on lines 39/40 in main_background.js.

3. Update the version number in manifest.json.

4. Add user-visible changes to NEWS.

5. Update the version number in docs/version.texi.

6. Commit and tag the release, and push to Savannah.

7. Run build.sh in a clean repo and upload the generated librejs.xpi to Mozilla.

8. Fix any validation warnings or errors during the upload in the previous step, test, re-tag the release, re-push to Savannah, and re-upload to Mozilla.

9. Generate docs from docs/ using gnulib gendocs.sh script (https://git.savannah.

   gnu.org/gitweb/?p=gnulib.git;a=blob;f=build-aux/gendocs.sh), and commit

the docs to the website cvs repo under the manual/ directory. This will update the

   docs at https://www.gnu.org/software/librejs/manual/.

10. From this step onwards it is required that Mozilla has approved the new version. Download the signed xpi from Mozilla website (https://addons.mozilla.org/en-US/

   firefox/addon/librejs/).

11. Follow the steps in Information for GNU Maintainer to upload the source tar ball (downloadable from cgit) and the signed xpi to the GNU FTP server.

12. Update the download links on the LibreJS webpage (https://gnu.org/s/librejs) to point to the new release.

13. Announce to info-gnu, help-librejs, bug-librejs mailing lists, and the Savannah project News section.

14. (Optional) Update projects that uses LibreJS like gnuzilla and abrowser.

Appendix A Installation Requirements

A.1 Mozilla Browser

You will need one of the many flavors of the Mozilla browser to use LibreJS. It can be installed on the following:

GNU IceCat, Mozilla Firefox, Trisquel Abrowser, Debian Iceweasel.

LibreJS works on these browsers starting from version 60. We recommend that you use the latest version of your Mozilla browser. LibreJS has been tested on a GNU/Linux distribution, but it is compatible any operating system as long as you’re using a compatible Mozilla browser.

Appendix B LibreJS Internals

LibreJS intercepts HTTP responses and rewrites their contents after analyzing JavaScript within them. It does not remove script nodes and attributes from the page, but instead “deactivates” them by replacing their content with a commented notice.

LibreJS detects the most common cases using the HTTP response method described above, but in less common edge cases, or when running code locally, LibreJS cannot detect JavaScript during the response stage.

Appendix C GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

 http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public.

A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition.

The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled ``GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.